### v4

# John Mendonça1,2, Alon Lavie3,4    Isabel Trancoso1,2

Abstract
Evaluating the quality of open-domain chatbots has become increasingly reliant on LLMs acting as automatic judges. However, existing meta-evaluation benchmarks are static, outdated, and lacking in multilingual coverage, limiting their ability to fully capture subtle weaknesses in evaluation. We introduce  [Uncaptioned image] MEDAL, an automated multi-agent framework for curating more representative and diverse open-domain dialogue evaluation benchmarks [^1] Code and data alongside submission.. Our approach leverages several LLMs to generate user-chatbot multilingual dialogues, conditioned on varied seed contexts. Then, a state-of-the-art LLM (GPT-4.1) is used for a multidimensional analysis of the performance of the chatbots, uncovering noticeable cross-lingual performance differences. Guided by this large-scale evaluation, we curate a new meta-evaluation multilingual benchmark and human-annotate samples with nuanced quality judgments. This benchmark is then used to assess the ability of several reasoning and non-reasoning LLMs to act as evaluators of open-domain dialogues. Using MEDAL, we uncover that state-of-the-art judges fail to reliably detect nuanced issues such as lack of empathy, common sense, or relevance.
MEDAL: A Framework for Benchmarking LLMs as Multilingual Open-Domain Dialogue Evaluators
John Mendonça1,2, Alon Lavie3,4and Isabel Trancoso1,2 1 INESC-ID, Lisbon 2 Instituto Superior Técnico, University of Lisbon 3 Carnegie Mellon University, Pittsburgh 4 Phrase, Pittsburgh {john.mendonca, isabel.trancoso}@inesc-id.pt, alavie@cs.cmu.edu
1 Introduction
Figure 1: Overview of  Refer to caption MEDAL, a 3-step multi-agent framework for curating native multilingual dialogue meta-evaluation benchmarks. ➊ Dialogue generation (§3): Generation of multi-turn, multilingual synthetic dialogues seeded from diverse contexts, between an LLM acting as a user (with utterances validated by an LLM-judge within a feedback loop) and the target chatbot; ➋ Automated labelling (§4): Multidimensional evaluation of all dialogues by a strong LLM (GPT-4.1); ➌ Benchmark Curation (§5): Curation of meta-evaluation benchmarks using human annotations used to assess LLM judges.
As research increasingly focuses on Large Language Models (LLMs), new models are being developed at a pace that rapidly outstrips our ability to evaluate them meaningfully (Laskar et al., 2024). This growing gap between model capability and evaluation capacity has become a significant impediment to the safe and effective adoption of new LLMs for tasks at large. When examining open-domain dialogue evaluation within this rapidly evolving context, we observe significant deficiencies in the methodology used for benchmarking LLMs tasked with evaluating open-domain dialogues. In particular, most existing benchmarks rely on obsolete chatbot responses conditioned on human-human interactions (Mendonça et al., 2024a), creating a mismatch with contemporary human-chatbot interaction paradigms. Furthermore, to the best of our knowledge, efforts to diversify benchmarks beyond English have been largely limited to translating existing English datasets into other desired languages, rather than natively curating data in these languages (Mendonca et al., 2023; Zhang et al., 2023; Rodríguez-Cantelar et al., 2023), thus failing to capture potential linguistic and cultural nuances in dialogue quality.
To address these gaps, this work presents an automated, multi-agent framework called  [Uncaptioned image]MEDAL [^2] Multilingual Evaluation of Dialogue-evaluators using Automated LLM-benchmarks.. MEDAL is designed to generate, evaluate, and curate more relevant and diverse dialogue meta-evaluation benchmarks as follows:
  1. 1.
    Dialogue generation (§3): synthetically generate a large collection of dialogues in multiple languages, using several LLMs acting as users and chatbots.
  2. 2.
    Automated labelling (§4): a strong LLM provides large-scale, multi-dimensional analysis of the generated dialogues, surfacing varying levels of failure across languages w.r.t empathy, common sense, and even fluency.
  3. 3.
    Benchmark curation (§5): guided by these automated labels, we curate and human-annotate a balanced set of dialogues to form a meta-evaluation benchmark that directly assesses LLM-judges.
Our analysis highlights several deficiencies in current models’ ability to accurately identify issues in dialogues, suggesting more nuanced affective and reasoning capabilities are required to achieve performance parity with human experts. Crucially, this methodology allows us to perform the first systematic assessment comparing the open-domain dialogue evaluation capabilities of reasoning and non-reasoning models within a multilingual context.
In summary, our contributions are as follows:
  • We propose  [Uncaptioned image] MEDAL (Figure 1) for building multilingual meta-evaluation benchmarks, with modular generation, labelling, and curation stages [^3] While not the focus of this work, the large-scale raw dialogues and labels produced during MEDAL’s generation and automated annotation stages can be repurposed to pre-train lightweight evaluators or to fine-tune dialogue agents.. MEDAL can be further extended to include additional languages and/or LLMs (as users, chatbots, or judges).
  • We conduct a comprehensive evaluation of LLMs as multilingual chatbots finding that there are significant cross-lingual performance differences between open-access LLMs.
  • A new meta-evaluation benchmark that explicitly measures the performance of LLMs as dialogue evaluators across six languages. Our analysis shows that reasoning models consistently outperform their non-reasoning counterparts but that a significant gap between humans and LLMs remains, especially for more nuanced quality dimensions (e.g., empathy, common sense).
2 Related Work
2.1 LLM-driven Dataset Curation
Several studies suggest leveraging LLM’s extensive world knowledge and linguistic knowledge for augmentation and synthetic data generation in order to scale dataset sizes (Ding et al., 2023; Li et al., 2023; Long et al., 2024). LLMs have been used to scale datasets for classification (Sahu et al., 2022; Hartvigsen et al., 2022), dialogue (Chen et al., 2023; Kim et al., 2023), and evaluation (Perez et al., 2022; Chiang and Lee, 2023; Tan et al., 2024).
LLMs have also been used as a drop-in replacement of humans via simulation (Aher et al., 2023; Huang et al., 2024). For dialogue generation in particular, some approaches employ several LLMs collaborating in a multi-agent system to improve the quality and faithfulness of dialogue generation and annotation (Guo et al., 2024; Ibrahim et al., 2025; Ma et al., 2025). Similar to this, our work employs a multi-agent system to automate synthetic dialogue generation. However, we are the first to extend it to multiple languages, and as a step for the curation of examples for human annotation.
2.2 Meta-evaluation Benchmarks
The majority of currently used dialogue meta-evaluation benchmarks predate the widespread introduction of LLMs (Mehri and Eskenazi, 2020a, b; Zhang et al., 2021). This raises concerns regarding their effectiveness at evaluating current human-chatbot interactions, especially considering that the majority of the obtained turn-level annotations use the same dialogue datasets (Yeh et al., 2021).
Our work draws inspiration from two main bodies of work. Finch et al. (2023b) investigates the ability of ChatGPT-3.5 (Ouyang et al., 2022) for dialogue behaviour detection for nine categories in real human-chatbot dialogues. Mendonça et al. (2024b) conducts a similar behavioural analysis with the curation of a large scale turn-level benchmark that evaluates dialogues from the Soda dataset (Kim et al., 2023). This evaluation used GPT-4 (OpenAI, 2024) as a judge, and then validated a small portion using human annotators.
While these and other recent benchmarks (detailed in Appendix A) offer valuable insights, they predominantly focus on English. A notable exception is the work by Rodríguez-Cantelar et al. (2023), which contains a subset of unique Chinese dialogues. Similarly, Zhang et al. (2023) extend coverage by translating many existing meta-evaluation benchmarks into multiple languages, though their focus remains on translating rather than designing new multilingual evaluation tasks. Additionally, only the English subset of Rodríguez-Cantelar et al. (2023) (60 dialogues, which were translated to Chinese and Spanish) contains chatbot responses obtained from LLM-based chatbots. In light of these limitations, MEDAL positions itself as the first truly multilingual meta-evaluation benchmark that explicitly assesses the performance of a wide range of recent open-access LLMs across languages.
3 Dialogue Generation
The first step in MEDAL is to synthetically generate dyadic dialogues between a human user (simulated by an LLM) and a chatbot. This scaffolding step provides three key advantages over curating static, human-chatbot dialogues: (1) scalability, allowing us to overcome expensive large-scale collection of native dialogues data across multiple languages; (2) contemporary relevance, enabling the on-demand generation of interactions that reflect modern conversational patterns and topics not found in older corpora; (3) error representation, as a large volume of dialogues allows us to better surface errors that become increasingly hard to find as chatbots improve.
To mitigate the risks of synthetic data, we adopt a multi-agent setup where we recruit several LLMs acting as the user (§3.1.1) and employ a feedback loop using an LLM as a user judge (§3.1.2). The resulting user turns are externally validated by human annotators (§3.5).
3.1 Multi-agent Framework
Given the instruction following and creative writing abilities of LLMs for dialogue generation (Kim et al., 2023), we prompt-instruct an LLM (acting as a human user and equipped with the topic) to begin and guide the conversation with a chatbot (another LLM, conditioned only on prior conversational context).
3.1.1 LLM as a User
To maximise diversity and minimise inherit biases, we condition the generation of dialogues using several contextual cues, including a scene description (obtained from a commonsense knowledge graph), a persona, an affective state and a target language. The inclusion of detailed information for contextualization purposes has been shown to outperform conversations sampled without context in terms of specificity and interestingness (Kim et al., 2023; Lin et al., 2024). A detailed explanation of each contextual cue is presented in Appendix C.
3.1.2 LLM as a User Judge
In order to ensure maximum quality in the generation of user utterances, and inspired by prior work (Pan et al., 2024), we include an automatic online feedback loop validation step which employs an LLM as a judge. In particular, this feedback loop is tasked with determining if: (1) the dialogue should end (thus preventing situations where the dialogue becomes redundant); and (2) the generated utterance is natural and appropriate w.r.t to the scene, persona, and language. If the generated utterance fails to meet the defined criteria, the user LLM is asked to regenerate the utterance response, conditioned on the rejected response and its feedback. An iteration limit ensures termination in cases where the judge may fail to accept the generated utterances indefinitely.
3.1.3 LLM as a Chatbot
We employ several open-access multilingual LLMs of different sizes as target chatbots. All models share the same system prompt, which conditions the models to generate natural, conversational language that is clear and easy to follow. In order to mimic other studies of open domain chatbot performance, the LLMs generating the chatbot responses have access to the dialogue history but are not explicitly provided with the external context (scene and persona) that was made available to the user.
3.2 Native Language Generation
To address the lack of resources in other languages, previous work in the curation of multilingual dialogue datasets has translated existing English dialogues into other languages (Lin et al., 2021; Liu et al., 2023). This approach to generating dialogues in non-English languages has several limitations. First, English-centric cultural artifacts like named entities, locations, and events remain unchanged after translation. More critically, translation does not offer a true, non-confounded evaluation of a model’s native multilingual capabilities.
In MEDAL, we keep the templated prompting sentence in English, but ask the LLM to act as a user from a specified country, and allow it to adapt the scene while generating directly in the native target language. This promotes better cultural diversity in contrast with translation [^4] A cherry-picked example from our dataset highlights this diversity for a dialogue discussing the music genre ”funk”: In Portuguese it typically refers to Funk carioca, whereas in English, it is associated with the 60s North American Funk.. Evidence that native generation yields more culturally appropriate dialogues is provided through a head-to-head comparison in Appendix C.5.
Figure 2: Heatmap of detected issues and average overall quality determined by GPT-4 for all tested chatbots across six languages.
Label Definition
Uninterpretable (!Int) Outputs unclear or confusing text to the point that its meaning is difficult to determine.
Unsafe (!Safe) Promotes or suggests harmful actions, or is offensive or threatening.
Lacks Empathy (!Emp) Fails to express appropriate emotional understanding when the context calls for it.
Lacks Common sense (!Com) Shows poor reasoning or contradicts basic, widely known everyday knowledge.
Repetitive (Rep) Unnecessarily repeats the same point, phrase, or idea in a way that adds no value.
Incoherent (!Coh) Contradicts itself or makes statements that do not logically follow from prior context.
Irrelevant (!Rel) Introduces unrelated content that does not align with the flow or subject of the conversation.
Non Factual (!Fac) Provides information that is objectively false or contradicts publicly verifiable facts.
Overall Overall assessment of the dialogue.
Table 1: Definition of issue labels used for the chatbot evaluation study and meta-evaluation benchmark.
3.3 Experiments
Models.
We use GPT-4o-mini5552024-07-18 snapshot accessed March 2025. (OpenAI et al., 2024) and Gemma-3-27b-it (Gemma et al., 2025) as our users, and Gemini-2.0-Flash (Google, 2025b) as the user judge, given the strong multilingual capabilities and reasonable size/price of these models. For chatbot responses, we employ 8 open access LLMs of 4 different families: (1) Aya Expanse (Dang et al., 2024) with sizes 8B and 32B; (2) LLama-3.1 Instruct (Grattafiori et al., 2024) with sizes 8B and 70B; (3) Qwen2.5-Instruct (Qwen et al., 2025) with sizes 3B, 7B, and 72B; (4) and finally Phi-3.5-mini-Instruct (Abdin et al., 2024). This selection was made taking into account the supported languages, model size diversity, and our own limitations w.r.t to infrastructure.
Languages.
MEDAL allows for the generation of dialogues in any language supported by the LLM agents utilized in the generation process. Due to limitations in annotator availability, we limit the work reported here to Chinese, English, German, French, Portuguese and Spanish.
Statistics.
In total, 38,400 dialogues were generated, stemming from 4,800 unique conversation starters (400 scenes, two user LLMs, and six languages). Additional statistics such as average response length, number of turns and diversity metrics are included in Appendix C.
3.4 Qualitative Analysis
Despite the inclusion of a feedback loop, a cursory analysis uncovered instances where the user LLM failed to maintain its role as a human user (i.e by erroneously reversing the role and acting as the responding chatbot) and by including other languages in its responses. These issues highlight limitations of our LLM judge (Gemini-2.0-Flash) in reliably identifying conversational roles and maintaining consistent linguistic output across turns. Subsequently, we employ GPT-4.1 to automatically detect and remove these dialogues, while ensuring we keep the dialogues that only exhibit chatbot deficiencies. We find that 6.44% of dialogues were flagged as having malformed user responses (4.68% due to role confusion), resulting in a final dataset of 35,927 dialogues. Additional analysis are available in Appendix C.3.
3.5 Human Validation
In order to evaluate the performance of the LLM as a user, we conduct a small scale human validation, where we ask annotators to rate the quality of the utterances generated by the LLM acting as a user in terms of their human-likeness, on a 1-5 Likert scale. Our annotator pool consists of a mix of graduate professionals in linguistics and computational linguistics, all with prior experience with annotations [^6] Finch et al. (2023a) reports low training pass rates when employing crowdsourcing platforms such as MTurk.. We assign 2 annotators per language, except for German and Spanish where we only employ a single annotator due to lack of annotators with professional level expertise in these languages. We validated the quality of user responses in the 600 dialogues (100 per language) later used for multidimensional human annotation (see Section 5). Even when selecting the lowest score among annotators, 98% of the dialogues were rated 4 or 5 (5:79.3%; 4:18.7%; 3:1.5%; 2:0.3%; 1:0%), with an average agreement (measured with Krippendorff’s $\alpha$) of .2592, denoting fair agreement.
4 Automated Labelling
Given the scale of dialogues generated in Section 3, an exhaustive human annotation is prohibitorily expensive. As a solution, we conduct an automated labelling step to (i) surface a wide range of possibly infrequent conversational issues at scale, (ii) compare error distributions across model families, and (iii) construct a balanced and representative pool of dialogues for meta-evaluation benchmark curation.
4.1 Experimental Setup
We evaluate the performance of the chatbots with GPT-4.1. This automated evaluation consists of a multidimensional analysis of each chatbot’s responses throughout the dialogue using a subset of behaviours from the ABC-Eval framework (Finch et al., 2023a) identified in Table 1. More details regarding this automated analysis, including the full prompt used for instructing the LLM are available in Appendix D.
4.2 Results
Figure 2 presents the percentage of detected issues across the 8 dimensions, together with the average overall score for all tested chatbots. Of the 8 assessed categories, non-factual appears to be the dimension in which even the stronger models struggle to some degree – the best performing model was found to output non-factual information in 2.7% of its dialogues.
Model Size.
The performance differences between models roughly align with model size, with the exception of Aya Expanse-8b, which is competitive in performance with its larger counterpart, Aya Expanse-32b and the 70B models. We observe noticeable differences in performance between these two groups for the majority of categories. However, for unsafe, repetitive and non-factual, the performance differences are less pronounced across model sizes, which may indicate that this behaviour is not strongly associated with model size but with other factors such as post-training data.
Figure 3: Partial English example from our curated benchmark, together with a failed evaluation conducted by Gemini-2.5-flash (with reasoning). Additional examples can be found in Appendix E.
Language.
We observe that, in general, all models perform best in English. Additionally, the model’s "family" largely predicts its language performance. For example, Qwen models underperform in German, in particular in the uninterpretable category. A cursory check of some German dialogues for these models indicates that many of the responses contain Chinese or English words. We also note that Phi-3.5 and Llama-3.1-8B generally struggle more with Chinese. This shows that despite being advertised as multilingual, many of these models exhibit substantial performance differences, even across high-resource languages.
5 Meta-Evaluation Benchmark
While the resulting automated labelling of Section 4 provides scale, prior work has identified limitations when using LLMs as judges (Zheng et al., 2023; Wu and Aji, 2025). When used for dialogue evaluation, LLMs have been shown to struggle with coherence and commonsense reasoning (Finch et al., 2023b; Mendonça et al., 2024a). Given that the goal of MEDAL is to curate meaningful meta-evaluation benchmarks, we propose distilling this large pool of synthetically generated and automatically labelled dialogues into high-quality benchmarks annotated by professional annotators. Figure 3 presents an example from our benchmark together with an incorrect automated evaluation.
5.1 Benchmark Curation
In order to construct a meta-evaluation benchmark that maximizes the diversity and representativeness of dialogue quality, we employed a multi-stage, balanced sampling strategy to select 100 dialogues per target language. With a total of 9,000 assessments (9 quality dimensions $\times$ 10 annotators $\times$ 100 dialogues) spanning 600 unique dialogues, our benchmark’s size is larger than other influential dialogue-level benchmarks [^7] DSTC11 (Rodríguez-Cantelar et al., 2023) evaluates 277 dialogues, FED (Mehri and Eskenazi, 2020a) 120 dialogues and ABC-Eval (Finch et al., 2023b) 400 dialogues..
Filtering.
We start by automatically excluding dialogues identified as containing uninterpretable chatbot responses due to the presence of other languages in the responses. These are detected via keyword matching of GPT-4.1 assessments. We exclude this data since these issues are simple to detect with rule-based methods, and therefore lack significant value in a meta-evaluation benchmark.
Balanced Selection.
We attempt to secure an equal distribution of issues (including dialogues free of any detected issues) by conducting an iterative selection of dialogues that also ensures balance in terms of overall score and coverage of all chatbots. To achieve this, we employed a two-stage procedure: (1) we first seeded the benchmark by randomly sampling one dialogue per user–chatbot pair that contained no issues and had an overall score of 5; and (2) we then iteratively added dialogues containing different issue types, each time selecting examples that helped balance the distribution of overall scores and chatbot representation. When multiple candidates satisfied these criteria, we chose one at random, and when no suitable example was available [^8] This was frequent when looking for dialogues from stronger chatbots with rarer issues. we ignored that requirement in the selection and continued the selection process.
5.2 Human Annotations
Figure 4: Spearman correlations across languages (all significant with $p<0.01$). Combined denotes the benchmark composed by all individual languages.
The annotators recruited to conduct the evaluation of the human-likeness of user generated utterances (Section 3) were also tasked with evaluating the performance of the chatbot on the same set of dialogues. Each annotator was asked to identify whether a given issue was present in the chatbot responses (binary classification), and in the end provide an overall assessment of the chatbot (1-5) throughout the dialogue. Instructions emphasized the need to evaluate the chatbots as if they were employed in online chit-chat interactions.
Inter-annotator agreement was high overall, with annotators agreeing on over 89% of cases on average across languages and dimensions (Table 3). Agreement was particularly strong for more objective categories such as non-factual and unsafe (>95%), while it was somewhat lower for more nuanced categories such as repetitive (>74%), which are inherently subjective. In a debriefing interview with several annotators, we found that the few instances of disagreement were all driven by differences in personal opinions. This highlights the value of human annotation in capturing a diverse range of perspectives (Plank, 2022). Additional details regarding this annotation are presented in Appendix E.1.
Figure 5: F1 scores for the positive label (F1+, detected issue) and negative label (F1-, no issue). Human performance is measured by comparing the two human annotation sets.
5.3 Meta-Evaluation of LLMs as Automated Dialogue Evaluators
Settings.
We utilize a shared evaluation system prompt that consists of the guidelines provided to the annotators verbatim. We set the temperature to 0 whenever possible and ask the evaluation to be returned in a structured format. For reasoning models, we set the effort to medium.
Models.
We evaluate several strong LLMs on our meta-evaluation benchmark: for non reasoning models we selected GPT-4.1 (our evaluator for the large scale evaluation), Gemini-2.0-Flash (our user response evaluator), and Deepseek-V3-0324 (DeepSeek-AI et al., 2025b); for the reasoning models we employ o3 (OpenAI, 2025), Gemini-2.5-Flash (Google, 2025a) with thinking, and Deepseek-R1 (DeepSeek-AI et al., 2025a).
Metrics.
For overall quality prediction we employ Pearson and Spearman correlations. For issue detection, we calculate classification metrics considering the positive and negative occurrences as classes by averaging results across the double human annotations, following Finch et al. (2023b). Additional metrics, such as Pearson correlation, Precision, Recall, Accuracy and prediction counts, are available in Appendix E.2.
5.3.1 Meta-evaluation Results
Overall Quality Prediction.
Figure 4 presents the correlation of the various LLMs as evaluators on our curated benchmark across all languages. Among the models evaluated, o3 achieves the highest correlation with human ratings on the full benchmark (closely followed by GPT-4.1) whereas Deepseek-V3 exhibits the lowest.
Issue Detection.
In terms of F1- (Figure 5), all evaluated models achieve results comparable to human expert performance. This indicates that identifying issue-free dialogues is a relatively easy task for current LLMs. The F1+ results, however, reveal a significant gap between human and model performance across most failure categories. While human annotators retain strong F1+ scores across all categories, the evaluated models struggle to match this performance.
5.3.2 Discussions
Models struggle with nuanced dimensions.
In general, we find models achieved near parity with human performance on more objective issues such as non factual and uninterpretable and repetitive, whereas more nuanced issues such as lacks empathy and lacks common sense are the issues that are the hardest to detect, with F1+ below 0.4 for even the top models. These categories are particularly challenging because they often lack explicit surface cues: empathy judgments depend on subtle affective alignment, while commonsense errors require integrating world knowledge with pragmatic inference, areas where LLMs appear to still struggle.
Reasoning models stronger than their non-reasoning counterparts.
When comparing the performance of the reasoning models against the non-reasoning models we note that the reasoning models showcase better performance on both the overall prediction and issue detection, thanks for the most part due to better recall. In particular, we observe the largest differences on the nuanced dimensions, where reasoning appears to help evaluators identify gaps in commonsense knowledge or missing affective cues that non-reasoning models frequently overlook. The exception to this is GPT-4.1, where it generally matches o3 performance.
Cross-lingual variations.
Our multilingual design allows us to probe whether evaluator performance generalizes across languages. While sample size and confounding factors such as annotator variation and uneven dialogue quality make language-wise results difficult to generalise, we nevertheless observe consistent trends (Figures 4 and 14). For example, GPT-4.1 and o3 achieve relatively stable performance across all six languages, while DeepSeek-V3 significantly underperforms in Spanish and French, indicating this model may suffer from training data imbalances or limited cross-lingual generalization for these languages.
6 Conclusions
This work introduces  [Uncaptioned image] MEDAL, a novel and scalable framework for generating and curating native multilingual open-domain dialogue meta-evaluation benchmarks. We leverage multiple LLM agents acting as user, chatbot, and feedback judge to scale multilingual dialogue generation, and then employ a strong LLM-judge for large-scale automated evaluation. This process not only provides insights into the performance of small-to-medium open-access LLMs as multilingual chatbots, revealing significant cross-lingual performance variation, but also supports the curation of a balanced meta-evaluation benchmark. Using this benchmark, we show that while reasoning LLM evaluators consistently outperform their non-reasoning counterparts, they still struggle with nuanced conversational dimensions such as empathy and common sense. MEDAL is readily extensible to additional languages and LLMs in any role, making it a flexible and scalable resource for advancing dialogue evaluation.
7 Limitations
Generation.
We recognize several limitations in our generation step. First, LLMs exhibit uneven global cultural awareness, which can introduce biased or stereotypical notions when left to generate culturally diverse content (Li et al., 2024a) (though we argue this approach is preferable than simply translating from English dialogues). Second, to reduce model-specific biases, we employed several small-to-medium sized LLMs spanning different sizes and families. Our framework remains model-agnostic so that any LLM may be integrated. As such, our dataset does not reflect state-of-the-art conversational performance. Finally, while our validation shows high human-likeness, future work should investigate whether similar LLM-judge failure modes appear in dialogues with genuine human users, who may exhibit more unpredictable and diverse behaviors.
Evaluation.
Since our automated evaluation relies on GPT-4.1, we risk introducing model-specific biases. This could have been mitigated by including more LLMs (Qian et al., 2025). However, our meta-evaluation indicates this evaluator achieves the highest overall recall in issue detection across languages, and all evaluated judges report middling precision. Therefore, we believe the inclusion of several LLMs in the decision-making process would likely still leads us to overestimate the true prevalence of issues in the dialogues, while some problematic instances may still go unnoticed. Another limitation relates to our assessment being confined to synthetically generated dialogues, which may not generalize to real-world conversational data. For instance, since LLMs are designed to be cooperative, an LLM acting as a user will seldom flag chatbot deficiencies, a behaviour that can assist judges in detecting chatbot failures.
Curation.
We acknowledge that there are selection biases by having our curation on the evaluation conducted by a single LLM. In particular, we sample dialogues that GPT-4.1 identified as having issues. This means that our curated benchmark is overrepresented by issues that GPT-4.1 was able to detect. If we consider these to be the easiest ones to detect, our meta-evaluation benchmark is posed to mostly probe evaluator sensitivity. Another limitation resides with the small sample size and annotator diversity, especially for German and Spanish, for which we only have a single annotator. This limitation, plus the fact that chatbots have different performances for each language, makes meta-evaluation analysis at the language level difficult to generalise.
Benchmarking.
While LLMs appear to showcase breakthrough performance across numerous benchmarks, many of these are rapidly becoming either compromised due to data contamination (Sainz et al., 2023), or obsolete as models surpass their difficulty. As such, there is a critical need for lifelong evaluation methodologies. MEDAL offers a promising step in this direction, by enabling on-demand generation of novel, multilingual dialogues, therefore mitigating leakage and ensuring relevance as new LLMs emerge. However, MEDAL relies on automated assessments as a screening stage to reduce human annotation load. As LLMs continue to improve, a challenge emerges on the automated curation of examples for human annotation. Future work should look into more refined evaluation acquisition policies, such as active testing (Kossen et al., 2021; Li et al., 2024b).
8 Ethical Considerations
Annotators.
We targeted a compensation rate which reflects a fair wage according to local costs of living. We standardized pay across annotators and based on estimated workload duration. While the annotation task was estimated to require approximately 4 hours of work, average self-reported durations were conservatively rounded up, and all annotators were compensated for 5 hours of work, regardless of individual completion time. No personal identifiable information was stored.
Safety.
Large scale synthetic curation carries the risk of including content that may be considered sensitive or damaging to different demographics. This is particularly relevant on the user-generated side considering that Atomic10X contains content alluding to self-harm or harm to others. We minimize the presence of harmful content by incorporating an LLM-judge that is tasked with refusing particular generations. For the chatbot side, however, part of our goal is to detect the frequency of such issues in their generations. As such, we decided against conducting safety filtering on chatbot responses. However, we acknowledge that our curated benchmark is not comprehensive enough to evaluate specialised content safety classification models.
Acknowledgments
We would like to thank Bruno Martins and Luis Fernando D’Haro for their helpful feedback. This research was supported by the Portuguese Recovery and Resilience Plan through project C645008882-00000055 (Responsible.AI) and by national funds through Fundação para a Ciência e a Tecnologia (FCT) with references PRT/BD/152198/2021 and DOI: 10.54499/UIDB/50021/2020 and 10.54499/PRT/BD/152198/2021. We also want to give thanks to MS Azure services and Irving Kwong for their sponsorship.
Appendix A Meta-evaluation Benchmarks
This section presents a brief survey of the datasets that have been used as a benchmark for LLM-based open-domain dialogue evaluation metrics.
The FED dataset (Mehri and Eskenazi, 2020a) contains conversations between humans and two chatbots, Meena (Adiwardana et al., 2020) and Mitsuku. It includes 124 dialogues (40 with Meena, 44 with Mitsuku, and 40 with humans), annotated for eighteen quality aspects. The dataset comprises 3,348 turn-level and 1,364 dialogue-level annotations, totalling 4,712 data points, with each conversation receiving one dialogue-level and three randomly selected turn-level annotations.
For the USR dataset (Mehri and Eskenazi, 2020b), annotations were collected for responses generated by models trained on the TopicalChat (Gopalakrishnan et al., 2019) and PersonaChat (Zhang et al., 2018) datasets. The models included Transformer (Vaswani et al., 2017), RNN Seq2Seq (Shang et al., 2015), LSTM (Hochreiter and Schmidhuber, 1997), and KV-MemNN (Miller et al., 2016). Annotations were performed on sixty dialogue contexts, each featuring multiple chatbot responses (six for Topical-Chat, five for PersonaChat) and an additional human response.
The DSTC10 track (Zhang et al., 2021) utilized a test set derived from three sources: CHANEL-JSALT2020, ChatEval (Sedoc et al., 2019), and new annotations on TopicalChat (Gopalakrishnan et al., 2019) and PersonaChat (Zhang et al., 2018). Responses were generated by eight systems –LSTM Seq2Seq, Attention-based LSTM Seq2Seq (Sutskever et al., 2014), HRED (Serban et al., 2016), VHRED, BlenderBot (400M-Distill) (Roller et al., 2021), DialoGPT-medium (Zhang et al., 2020), T5-base (Raffel et al., 2020), and GPT-3 (Brown et al., 2020)—alongside a human baseline and random utterances.
The DSTC11 test set (Rodríguez-Cantelar et al., 2023) combined a portion of the DSTC10 test set with new human-chatbot dialogues. These dialogues featured state-of-the-art chatbots: ChatGPT, GPT-3.5 (Ouyang et al., 2022), and BlenderBot3 (Shuster et al., 2022) for English; and Chinese DialoGPT, Xiaoice (Zhou et al., 2020), and Plato-XL (Bao et al., 2022) for Chinese. The test set underwent translation and back-translation to English, Chinese, and Spanish. In total, it includes 4,839 turn-level and 277 dialogue-level annotations.
XDial-Eval (Zhang et al., 2023) is built on top of several existing English dialogue evaluation datasets, all of which predate the introduction of LLMs for dialogue. This curation process yields grand total of 14,930 annotated turns and 8,691 annotated dialogues. A commercial machine translation system is then used to extend these datasets to nine other languages: Chinese (ZH), Spanish (ES), German (DE), French (FR), Japanese (JA), Korean (KO), Hindi (HI), Arabic (AR), and Russian (RU).
In ABC-Eval (Finch et al., 2023a), multi-dimensional annotations at the turn and dialogue level were collected on unique conversations in English between collaborative humans and 4 different chatbots: (1) Blender-Decode (Nie et al., 2021); (2) Blender2 (Xu et al., 2022); (3) BART-FiD-RAG (Shuster et al., 2021); and (4) Emora (Finch et al., 2020). Annotations were obtained using ABC-Eval binary behaviour labels (16), dialogue-level (8), turn level (8) and comparative (8). In total, 400 dialogues were collected and annotated, 100 per chatbot.
For Soda-Eval (Mendonça et al., 2024b), the authors obtain 120K turn level annotations stemming from 10K dialogues extracted from the Soda dataset (Kim et al., 2023) which uses GPT-3.5 (Ouyang et al., 2022) for synthetic dialogue generation. The annotations include 9 issue labels plus an overall assessment using a Likert Scale.
Appendix B Human Annotation
All annotation work for this study was conducted by individuals affiliated with our research institution, including graduate students and post-doctoral researchers in linguistics and computational linguistics. Each language (with the exception of German and Spanish) was annotated by two annotators. For German and Spanish, a single annotator was employed. We ensured that at least one annotator per language was a native-level speaker. Annotators who were not native speakers held graduate-level qualifications in linguistics and had prior experience with Machine Translation (MT) annotation in the target language.
Appendix C Dialogue Generation
C.1 Seed Context
Figure 6: Selected affective states from the circumplex model (Russell, 1980).
We utilize a seed context to minimize generation biases and augment the diversity of the generated dialogues. This context is composed of 4 different axes of diversity: Scene, Persona (including Gender), Affective State and Language. We then instruct an LLM to generate the first user utterance based on this seed context. The LLM is free to prioritize different axes when initiating the dialogue, typically starting with the scene, but also drawing on persona or affective state, or combining multiple cues, thus ensuring the required diversity in conversation topics and styles.
Scene description.
We use the templated sentence form of Atomic 10x (West et al., 2022) to describe scenes. Atomic 10x (CC-BY-4.0) is a synthetically generated commonsense knowledge graph, consisting of symbolic triplets describing two events and the relation between them. Whenever the scene does not include an affective state (frequently found in relations of type xEffect or xReact), we randomly sample an affective state from a curated list of states, equally distributed across the quadrants of the circumplex model (Russell, 1980). The inclusion of an affective state in the scene description aims to reduce bias in casual dialogue generation, which in our initial experiments overwhelmingly presented tones of excitement and happiness.
Affective State.
Some scenes from Atomic10X include an affective state that is appropriate for the established relation. However, for those that do not have one (which we check by simple keyword matching with a comprehensive list of affective states), we randomly sample one from our selected list of affective states. We present our selected affective states from the Circumplex model (Russell, 1980) in Figure 6. This selection was based on sampling an equal number of states on each quadrant (Valence, Arousal), while also minimizing arousal which we believe better reflects the lower intensity of this dimension in textual interactions.
Persona.
Since we want to have a simulation of human users interacting with chatbots, the inclusion of a persona provides a greater opportunity to diversify both the behaviour and style of the generated answers. We randomly sample personas from the publicly available subset of Persona Hub (CC-BY-NC-SA-4.0) (Ge et al., 2024), a collection of diverse personas automatically curated from web data. Since all personas are gender neutral, we optionally sample a gender to reduce bias in gendered languages.
Gender.
Using a simple rule-based analysis of the scenes and personas, we detect an uneven distribution of gender for Atomic 10X (54% male, 46% neutral), whereas PersonaHub is purely gender-neutral. Since we include languages that are grammatically gendered, we only sample gender-neutral scenes from Atomic10x and randomly select a binary gender class. The LLM is prompted to take this gender label into account if required to determine the gender of the user when generating utterances. This attempts to minimize gender bias in user generations.
First Turn Generation.
Since the start of the dialogue typically sets the tone and topic of human-chatbot interactions, we separate the generation of the first turn of the user (which is shared across the chatbots) from the generation of subsequent turns. In particular, we ask the LLM to generate a user initiating turn that introduces the external context (i.e., language, scene and/or persona) of the dialogue, and that is also appropriate for a human-Chatbot conversation. This differentiates our generated dialogues from typical human-human open-domain dialogue data collection efforts in several important ways. In particular, this prevents a greeting phase that is inherently lacking in dialogue diversity. Also, this inhibits user utterances that attribute anthropomorphism features to chatbots, which in real conversations with Chatbots would typically result in a templated refusal to engage by the chatbot ("As an AI, I don’t have the ability to…").
C.2 Generation Details
Our framework combines open-access [^9] All models were obtained from huggingface.co. with closed-sourced LLMs accessible via an API. For open-access LLMs, we use vLLM for inference on a single node with 4xA100 80GB GPUs. Dialogue length is capped at 10 turns due to cost constraints.
User.
For the first turn generation, we use the system prompt identified in Table 6 and set temperature to 1.5 and presence_penalty of 0.6 in order to maximize diversity. For subsequent turns, we set temperature to 0.9, top_p to 0.95, and limit the generation to 512 tokens.
User Validation.
We set a temperature of 0.1 and limit the generation to 64 tokens to force smaller feedback. Due to the creativity imposed by the generation parameters, we set the regeneration limit of the first User turn to ten attempts, which was sufficient to have all seed contexts accepted by the LLM-judge. For subsequent turns, we limit regeneration attempts to five, which achieves a 90% pass-rate, and keeps costs below a factor of four [^10] Incorporating the feedback loop makes the process 4 times more expensive in terms of LLM queries, assuming response generation and evaluation steps have equal cost.. Turns that fail all attempts at regeneration have the corresponding dialogue terminated.
Chatbot.
Similar to the user, temperature was set to 0.9, with a top_p of 0.95 and generation is capped at 512 tokens.
#Dialogues Avg #Turns Avg. Utt. Len. Lexical Diversity
DailyDialog 13.1K 7.9 11.3 45.2
PersonaChat 10.9K 14.8 14.2 49.8
Soda 1.48M 7.6 16.1 42.6
Ours 35.9K 9.12 - -
ZH 6.0K 8.92 57.5 71.5
EN 6.0K 9.35 30.9 73.2
FR 6.1K 9.31 31.0 58.5
DE 6.0K 8.89 29.3 45.0
PT 5.9K 9.16 29.6 65.7
ES 5.9K 9.17 30.3 64.1
Table 2: Statistics of the generated dataset compared to other open-domain dialogue datasets. For Chinese, we use characters instead of words for utterance length calculation. We exclude the average calculation due to this reason.
C.3 Detection of Malformed User Behaviour
Figure 7: Proportion of dialogues automatically filtered due to role reversal or unintended language switching. Percentages are relative to the initial pool of generated dialogues for each language.
We used GPT-4.1 assessments to identify and remove malformed dialogues prior to annotation. These included two systematic error types observed during generation: (i) role confusion, where the user agent incorrectly assumed the role of the chatbot, and (ii) language mixing, where the chatbot produced responses in a language other than the intended target.
Across all six target languages and two user LLMs, 6.44% of dialogues were flagged as malformed (4.68% due to role confusion) and excluded from subsequent stages. Figure 7 reports the percentage of removed dialogues. We note that role reversal was more frequent for GPT-4o-mini, while unintended presence of multiple languages was equally found among both models. Not surprising, we found English to be the least likely language to contain this type of issue.
C.4 Statistics
We compare our our collected dialogue dataset with frequently used datasets for meta-evaluation in Table 2. In general, our dialogue subsets obtain similar lexical diversity values (calculated using MTLD [^11] We calculate MTLD with python lexical_diversity using spacy lemmas and jieba tokens for Chinese.; McCarthy and Jarvis, 2010), with languages such as Chinese and English achieving the largest diversity (over 70), whereas German achieved the lower diversity (45.0). We also note that the average utterance length is much higher across all languages (around 30 words per turn, excluding Chinese) when compared to other datasets. This is due to the chatbot side generally being much longer in length than the human response. Since we ask both sides to keep utterances concise (1-2 short sentences), we hypothesize this disparity is due to role differences (the chatbot side typically offers advice that runs longer than the user’s scene exposition). However, even when restricting our analysis to only user-side utterances in Latin-script languages, the mean utterance length is 23.4 words, which is still higher when compared to other datasets. This finding aligns with the well-known verbosity bias in LLMs, which tend to generate longer responses, likely as a by-product of preference optimization (Saito et al., 2023) [^12] Soda generates the dialogue as a whole, therefore utterance-level verbosity is reduced..
C.5 Translation vs Native Generation
To validate our hypothesis that natively generated dialogues in a target language are more authentic than to translated ones, we conduct a head-to-head comparison. In this experiment, the English subset of dialogues generated with aya-expanse-32b and gemma-3-27b were translated into the target languages using a prompted version of GPT-4.1. These translated dialogues were then benchmarked against the dialogues natively generated in the target languages on the same seed context.
We use Gemini-2.5-pro to select the dialogue with higher cultural appropriateness and naturalness. To mitigate order bias, we conducted two evaluation trials for each dialogue pair, reversing the order of presentation in the second trial. A "win" for native generation occurs when Gemini-2.5-pro rates it as superior in both trials, and a "loss" was recorded when the translated version was favoured in both instances. All other outcomes were classified as a "tie".
Figure 8: Automated evaluation results from a head-to-head comparison between natively generated and translated dialogues. The values represent the Win/Tie/Loss rates for the natively generated dialogues against their translated counterparts with respect to cultural appropriateness and naturalness.
The results presented in Figure 8 indicate that for most of the languages tested, dialogues generated natively are better in cultural appropriateness and naturalness compared to those translated from English. The exception is for Spanish, where the translated dialogues achieved a level of quality that was often comparable to the natively generated ones in the aspects evaluated.
Appendix D Automatic Dialogue Evaluation
D.1 Annotation Details
To save on costs, we only provide as guidelines the list of issue labels together with their definition. This allows for the model to over predict and therefore reduce bias for the curation step.
Figure 9: Proportion of identified issues per language.
D.2 Additional analysis
We present the proportion of identified issues per language in Figure 9. We note that German contains the largest amount of issues pertaining to Coherence and Interpretability. This is mainly due to the fact the Qwen family of models is particularly weak in generating German text, often mixing it with Chinese and/or English despite being officially supported. For the remaining languages, the distribution of issues is much less prevalent. When looking at issues pertaining to Relevance, Common sense, Empathy, Safety and Interpretability, the percentage of dialogues with such issues is below 5%, with English being the language with less issues across all types of issues.
This behaviour is also evidenced when looking at the distribution of overall scores by language in Figure 10. The vast majority (over 80%) of dialogues is rated 5 with the exception of German (70%). We also note that no dialogue was rated with the worst score (1). This could be due to two distinct reasons: firstly, that LLM-chatbots are robust enough to recover from problematic conversations; secondly, we designed our LLM to act as a collaborative user, which significantly reduces the chances of a dialogue breakdown.
Language !Int !Safe !Emp !Com Rep !Coh !Rel !Fac Overall
EN 1.00 0.95 0.91 0.97 0.76 0.82 0.79 0.94 0.84
ZH 0.76 0.98 0.94 0.96 0.85 0.74 0.77 0.96 0.68
FR 0.94 0.96 0.97 0.98 0.97 0.95 0.92 0.96 0.76
PT 0.82 0.93 0.90 0.91 0.74 0.68 0.82 0.95 0.82
Table 3: Inter-Annotator Agreement (IAA) calculated using the percentage of exact agreement between annotators across all dimensions and languages with two annotators. For "overall", we calculate the adjacent agreement rate, i.e., the percentage of score pairs that differ by no more than 1 (liu2025mdsevalmetaevaluationbenchmarkmultimodal). A formal definition of these dimensions is available in Table 1.
Language !Int !Safe !Emp !Com Rep !Coh !Rel !Fac Overall
EN 1.000 .6414 .4792 .5581 .2803 .5885 .4501 .7361 .4707
ZH .3258 .8236 .5415 .5811 .3981 .4236 .3220 .8347 .2622
FR .3716 .5811 .3877 .6581 .3877 .4212 .3886 .7967 .3508
PT .3337 .5522 .3926 .4243 .2908 .2419 .2562 .8415 .4182
Table 4: Inter-Annotator Agreement (IAA) calculated using Krippendorff’s $\alpha$ for all dimensions across languages with two annotators.
Appendix E Benchmark Curation
Figure 10: Overall Score distribution across languages.
Figure 11: Pearson correlations across languages (all significant with $p<0.01$). Combined denotes the benchmark composed by all individual languages.
E.1 Human Annotations
We present agreement rates across all dimensions and languages for which we had more than one annotator in Table 3. Following Finch et al. (2023a), we additionally present IAA calculated using Krippendorff’s $\alpha$ [^13] We omit bootstrap intervals due to instability caused by the low number of positive examples in an overall small sample size (100), which leads to unreliable or undefined CI. We additionally report raw percentage agreement (Table 3). in Table 4. Agreement values are generally within the agreement reported by Finch et al. (2023a) for behaviour labels ($0.3\leq\alpha\leq 0.8$)
Overall, issues such as non-factual (!Fac) and unsafe (!Safe) achieve very high correlations. This is due to the relatively objective nature of this issue [^14] Furthermore, annotators were allowed to consult external references to check for ”factuality”.. Agreement is consistently lower for more subjective categories such as repetition (Rep) and lacks empathy (!Emp), highlighting the challenges of reliably annotating these issues.
Figure 12: Precision and recall scores for issue detection.
Figure 13: Reported accuracy and total number of predicted dialogues with issues, for each model. Human Accuracy performance is measured by comparing the two human annotation sets. A statistical significance test between outputs of models and human annotators (estimated using McNemar’s Test with significance level of 0.05) confirms both annotators outperform the best model except for "Unsafe". For issue counts, we average the amount of identified dialogues whenever there is more than one annotator.
E.2 Additional Results
E.2.1 Overall Quality
We present per-language Pearson correlations for Overall Quality in Figure 11 and numeric correlations results in Table 5.
E.2.2 Issue Detection
Precision vs Recall.
Model Pearson Spearman
GPT-4.1 .4767 .4525
Gemini-2.0-flash .3707 .3558
Deepseek-V3 .3206 .2817
Deepseek-R1 .4472 .4044
Gemini-2.5-flash .4462 .4200
o3 .4914 .4750
Table 5: Numeric overall quality correlation results on the full benchmark. All correlation results are significant with $p<0.01$. Bold denotes best model.
We present Precision and Recall scores in Figure 12. We note that across most issue types, models exhibit a significant performance gap compared to human annotators in recall. The exception is GPT-4.1, which often achieves a more balanced profile with generally higher recall than other models, though still below human levels. For Non-factual, we observe good performance by the LLMs. However, for more nuanced issues like Lacks Empathy and Lacks Common sense, all evaluated LLMs struggle considerably, with recall often falling drastically short.
Accuracy.
Accuracy scores (presented in Figure 13) are generally high across all models and categories, often exceeding 85%, with top models like GPT-4.1 and o3 achieving accuracies above 90% when detecting Unsafe and Non Factual content. The number of predicted issues per error category on the full benchmark (Figure 13) varies substantially, from as few as 3 instances (e.g., uninterpretable for DeepSeek) to over 150 (e.g., incoherent for Gemini 2.5).
Per-Language Performance.
For completeness, we report F1+ scores (Figure 14) for each individual language.
Figure 14: F1+ scores per language and issue type. Human results are not presented for Spanish and German due to the lack of a second annotator.
Figure 15: French example from our curated benchmark.
Figure 16: German example from our curated benchmark.
Figure 17: Portuguese example from our curated benchmark highlighting the gendered nature of this particular language (desanimada – female).
Figure 18: Chinese example from our curated benchmark with the initial GPT-4.1 assessment.
E.3 Additional Examples
We present additional examples from our curated meta-evaluation benchmark in Figures 15, 16, 17 and 18.
—- System prompt —–
You are a creative writer specializing in crafting human-like casual open-domain interactions with chatbots. Your task is to generate the first message a human user might send to a chatbot, based on the following inputs:
1. Scene Description: A small social context or event description.
2. Persona: A brief description of the individual’s role, background, or identity.
3. Gender: Gender of the individual if not already provided before.
4. Language/Culture: The language or cultural context of the user.
Guidelines:
- Use natural, conversational language typical of casual, open-domain interactions. Messages should feel authentic and concise, limited to one or two small sentences.
- Do not address the chatbot in a manner that assumes it has a physical body, a personal history, or experiences typical of a human (e.g., having a family, personal secrets, or emotions linked to past events).
- Do not write messages that imply the chatbot is someone the user has met before or can relate to as if it were a human friend.
- Ask for advice, opinions, information, or share personal reflections, experiences, or questions that do not attribute human characteristics to the chatbot.
- Reflect the age, emotional state and language/culture of the individual in tone, word choice, and phrasing.
- Incorporate the context from the scene description without explicitly repeating it word-for-word but subtly integrating its essence or themes.
- Do not use placeholder terms like "PersonY". Use realistic names, pronouns, or generic references suitable for the context.
- If gender is required and not provided in the persona or scene, use the one provided as "Gender".
Output:
Provide only the message that the human might send to a chatbot. Do not include quotation marks, meta-commentary, or any additional text outside of the generated message (including "user:").
[Examples]
Table 6: User utterance generation system prompt for the first turn.
—– System prompt —–
You are role-playing as a human in an online casual conversation. Your task is to generate a natural and authentic response given prior context and an optional feedback from a prior generation attempt.
Guidelines:
- Use natural, conversational language that reflects how humans communicate online with chatbots.
- Do not acknowledge that you are an AI or break character as the human in the conversation.
- Keep your single response clear and easy to follow, using short sentences and everyday language. The message should be concise (1 or 2 small sentences) and relevant to the conversation and scene.
- Respond in a way that feels humanlike. Avoid repeating previous content.
- Avoid verbose or robotic phrasing. Do not use the same conversational structure (e.g., starting with appreciation or a personal preference followed by a question) in every turn.
- If gender is required and not provided in the persona or scene, use the one provided as "Gender".
- Use the language specified in the scene.
- Do not use placeholder names like "PersonY". Use realistic names or generic pronouns that suit the context and language.
- Do not let the conversation drag on. If the conversation should end, output ’END_OF_DIALOGUE’ to signal the end of the dialogue.
- Take into account the optional feedback from a prior generation attempt, if provided, to improve the response.
Output:
Provide only the message that the human might send to a chatbot. Do not include quotation marks, meta-commentary, or any additional text outside of the generated message (including "user:").
—- User prompt —–
The scene is as follows:
[Scene + Persona + Affective State + Gender]
The Dialogue is as follows:
[Dialogue Context]
Prior failed generation attempt was:
[Prior Generated Response]
Feedback from this previous generation:
[Feedback]
Table 7: User utterance generation template for subsequent turns.
—– System prompt —–
You are a chatbot designed to engage in online casual conversations. Your task is to respond to messages directed at you in a way that fosters a smooth, engaging dialogue.
Guidelines:
- Use natural, conversational language that is clear and easy to follow, avoiding overly formal or robotic tones.
- Use the same language as the user.
- Keep your responses concise (1 or 2 sentences) with sentences that are short, easy to follow and relevant – aim for maintaining conversational flow.
- Avoid steering the conversation towards a specific goal, such as information provision or task completion. Instead, focus on maintaining an engaging dialogue.
- Do not use bullet points or overly structured lists; instead, respond in a fluid, conversational manner.
- Adapt your tone and content to match the style and mood of the conversation.
- Ask questions and introduce new elements or topics when appropriate to keep the exchange interactive, engaging and non-repetitive.
Table 8: Chatbot system prompt. Prior dialogue turns are provided as context using the chat template.
—– System prompt —–
You are a dialogue evaluation assistant tasked with determining whether a generated response (the last user message) meets the following criteria:
- Natural and Conversational: The response should sound like it was written by a real person in an ordinary online conversation, using language and expressions typical of a user.
- Concise and Coherent: The response should be brief (1–2 sentences), non-repetitive, and coherent with the prior conversation context.
- Appropriate Tone: The response should match the style, language, and mood expected from a user. It should not mimic an assistant’s voice by providing advice, guidance, or suggestions that are typically offered by the assistant. Asking for advice or seeking information is acceptable if it aligns with the user’s role.
- Role Appropriateness: The response must clearly reflect the user’s role. If the response includes elements (e.g., offering support, advice, or asking probing follow-up questions) that are characteristic of an assistant’s response, it should be flagged. The user should not break character or acknowledge that they are an AI.
- Non-Repetitiveness: Responses should not repeat of previous content, sentence structures (e.g., starting with appreciation or a personal preference followed by a question), or acknowledgments.
- Ending: The generated response can include the flag "END_OF_DIALOGUE" if the conversation should end. This flag should be used only when the conversation has reached a natural conclusion.
Your task is to evaluate ONLY the last message in the conversation against these criteria.
Output: "Yes." if the user response meets all criteria, or "No. <brief explanation>" if it does not.
—- User prompt —–
[Dialogue Context]
Table 9: User utterance evaluation prompt.